OpenStudy (anonymous):

Compute and write your answer in scientific notation. 8.2 x 105 2.75 x 10-3

OpenStudy (anonymous):

8.2/2.75=2.98181 105/10-3=108 2.981*108

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

are you multiplying or dividing?

OpenStudy (anonymous):

its dividing I thought?

OpenStudy (bloomlocke367):

okay... you didn't say in the question XD

OpenStudy (anonymous):

yes it has to be dividing

OpenStudy (bloomlocke367):

so you have\[\frac{ 8.2 \times 10^5 }{2.75 \times 10^{-3} }\] right?

OpenStudy (anonymous):

yes

OpenStudy (bloomlocke367):

okay... when you divide exponents, it's a little different.

OpenStudy (bloomlocke367):

you subtract instead of add... which I think you got XD so you're \(\Large\tt\color{lime}{CORRECT!!}\)

OpenStudy (anonymous):

awesome thanks! can you help me with one more for the night I've been struggling with this one

OpenStudy (bloomlocke367):

of course :) just be sure to open a new question XD